# suppSupplementary References

### v2

@@ Abstract
Human Mesh Recovery (HMR) from a single RGB image is a highly ambiguous problem, as similar 2D projections can correspond to multiple 3D interpretations. Nevertheless, most HMR methods overlook this ambiguity and make a single prediction without accounting for the associated uncertainty. A few approaches generate a distribution of human meshes, enabling the sampling of multiple predictions; however, none of them is competitive with the latest single-output model when making a single prediction. This work proposes a new approach based on masked generative modeling. By tokenizing the human pose and shape, we formulate the HMR task as generating a sequence of discrete tokens conditioned on an input image. We introduce MEGA, a MaskEd Generative Autoencoder trained to recover human meshes from images and partial human mesh token sequences. Given an image, our flexible generation scheme allows us to predict a single human mesh in deterministic mode or to generate multiple human meshes in stochastic mode. MEGA enables us to propose multiple outputs and to evaluate the uncertainty of the predictions. Experiments on in-the-wild benchmarks show that MEGA achieves state-of-the-art performance in deterministic and stochastic modes, outperforming single-output and multi-output approaches.
@@ 1 Introduction
Perceiving humans from images is a long-standing problem in computer vision, with applications in diverse fields such as sports [20, 78] or e-commerce [51, 85]. Many approaches rely on statistical body models like SMPL [52] for representing humans. Earlier human mesh recovery (HMR) methods recovered the SMPL pose and shape parameters from 2D cues using optimization-based techniques [6, 43]. However, these optimization procedures require good initialization, are time-consuming, and often converge to suboptimal minima. With the advancement of deep learning and the availability of datasets of images with 3D human pose and shape annotations, most approaches have shifted to a regression-based paradigm. Early approaches used architectures based on convolutional neural networks (CNNs) and multilayer perceptrons (MLPs) [34, 46]. Recent works have adopted Transformers [76] for extracting image features or making predictions [11, 24]. Despite achieving unprecedented accuracy, state-of-the-art HMR models still have some weaknesses, such as producing unrealistic predictions, especially when dealing with occlusions. To address these issues, recent works have proposed tokenizing the human pose using vector quantized variational autoencoders (VQ-VAEs) [21, 19]. This approach uses a discrete representation of the human mesh, learned from large-scale motion capture datasets, to confine predictions to the space of anthropomorphic meshes using a dictionary of valid mesh tokens. This tokenization aligns well with Transformer-based architectures, which were originally designed for processing discrete data in natural language processing. Notably, VQ-HPS [21] reframed HMR as a classification task and achieved state-of-the-art, demonstrating the great potential of human mesh tokenized representations in HMR.
While significant progress has been made in HMR, a major issue remains unaddressed in most prior works: a 2D image, especially with occlusions, cannot provide sufficient information to estimate a 3D human mesh with certainty [57, 71]. This limitation causes single-output models to be biased toward most common poses and body shapes [14]. To mitigate this problem, several works have proposed probabilistic approaches that generate multiple predictions from a single image [68, 41]. These approaches have used various families of generative models, ranging from conditional variational autoencoders (CVAEs) [70] to diffusion models [10]. However, this increase in diversity typically comes at the cost of accuracy [69], and none of these multi-output methods are competitive with the latest single-output HMR models.
@@
Figure 1: Human mesh recovery from a single image is an ill-posed problem due to depth ambiguity. Probabilistic approaches have aimed to address this by generating multiple predictions, but diversity often sacrifices accuracy. Introducing MEGA, our HMR model based on masked generative modeling achieves state-of-the-art performance on in-the-wild benchmarks while also quantifying uncertainty. Bluish regions in the mesh indicate low uncertainty, while reddish areas signify higher uncertainty.
@@
In this work, we introduce MEGA, a multi-output HMR approach based on self-supervised learning and masked generative modeling of tokenized human meshes. MEGA relies on the Mesh-VQ-VAE of [21] to encode/decode a 3D human mesh to/from a set of discrete tokens. Our training process unfolds in two steps: (1) Firstly, akin to (vector quantized) masked autoencoders [27, 3, 65, 64], we pre-train MEGA in a self-supervised manner to reconstruct human mesh tokens from partially visible inputs. This leverages large amounts of motion capture data without the need for paired image data. (2) Subsequently, for HMR from RGB images, we train MEGA to predict randomly masked human mesh tokens conditioned on image feature embeddings. During inference, we begin with a fully masked sequence of tokens and generate a human mesh conditioned on an input image. We propose two distinct generation modes: (2.a) In deterministic mode, MEGA predicts all tokens in a single forward pass, ensuring speed and accuracy; (2.b) In stochastic mode, the generation process involves iteratively sampling human mesh tokens, enabling MEGA to produce multiple predictions from a single image for uncertainty quantification.
@@
We evaluate MEGA on in-the-wild HMR benchmarks, comparing it to single-output and probabilistic HMR methods. MEGA achieves state-of-the-art (SOTA) performance when predicting a human mesh in a single forward pass in deterministic mode. In stochastic mode, MEGA outperforms both deterministic and probabilistic methods with a single prediction and significantly enhances its performance with an increased number of samples. This mode generates diverse, realistic human meshes, allowing for uncertainty evaluation and the proposal of multiple plausible outputs given an image. Additionally, we demonstrate that the pre-trained MEGA can generate realistic and diverse random meshes, further showcasing the flexibility of our approach.
In summary, we make the following two key contributions [^1] Code and trained models will be released upon acceptance.:
@@ 2.1 Human mesh recovery
Single output HMR. Since the release of the HMR [34] model, most approaches for recovering human meshes from images have been regression-based, using neural networks to make predictions directly from the image. These regression-based HMR methods can be categorized into parametric and non-parametric approaches. Parametric methods aim to recover the parameters of the SMPL model [72, 37, 45, 46, 24, 19, 39, 33, 83, 36]. They typically produce realistic predictions; however, some works have argued that the SMPL model parameter space is not the most suitable for predicting human meshes [14, 40, 12], leading to the development of non-parametric approaches. Non-parametric methods predict the coordinates of 3D vertices without relying on a learned parametric model. While earlier approaches used graph convolutional neural network architectures inspired by the mesh topology [40, 48], recent non-parametric models predominantly employ Transformers [11, 47, 18, 21, 53]. Although non-parametric methods yield accurate results, they can sometimes produce non-anthropomorphic meshes, particularly when training data is scarce [21].
In this work, we approach HMR differently by formulating it as a classification task. Instead of predicting SMPL model parameters or 3D coordinates, we aim to recover sequences of indices that can be decoded into a human mesh using the Mesh-VQ-VAE from [21].
Multi-output HMR. Estimating a 3D human mesh from a single image is challenging due to the depth ambiguity, especially when the person is partially occluded. Several works have proposed making multiple predictions to address the ill-posed nature of the problem. Earlier works employed compositional models [32] or mixture density networks [44]. More recent approaches rely on sophisticated probabilistic distributions [68, 67] and generative models, such as CVAE [70], normalizing flows [41, 4, 69], and diffusion models [10]. While these methods can predict diverse plausible solutions, they often face a trade-off between accuracy and diversity [69] and struggle to achieve SOTA results even when making multiple predictions.
@@ 2.2 Self-supervised learning for HMR
While prior works have demonstrated the importance of pre-training for the backbone of HMR models [5, 60], we propose pre-training the generative model on human meshes to leverage extensive motion capture data. This pre-training not only provides us with an unconditional human mesh generative model but also proves beneficial for training a robust HMR model. In deterministic mode, the encoder of our model can be discarded, as the decoder alone is sufficient for predicting sequences in a single forward pass (see Sec. 3.4). To our knowledge, this is the first work that discards the encoder of an MAE after pre-training; previous works typically discard the decoder and use the encoder to extract features.
@@ 2.3 Masked generative modeling
Masked modeling was introduced in BERT [16] for language modeling and extended to images with the MAE [27]. This technique trains a model to predict randomly masked tokens in a sequence based on visible tokens. Masked generative modeling builds on this by training a model to generate new samples, starting from a fully masked sequence and iteratively predicting a fixed number of tokens at each step [8, 7].
In this work, we develop a masked generative modeling approach for HMR. Using a tokenized representation of the human mesh is particularly well-suited for this task, as it allows for straightforward masking and replacement of mesh parts with mask tokens.
@@ 3 MEGA
3.1 Human mesh tokenization
@@ 3.2 Model
Overall pipeline. MEGA is a masked generative model based on an encoder-decoder Transformer architecture, illustrated in Fig. 2. During training, the Mesh-VQ-VAE encoder converts a 3D human mesh into a sequence of $N$ tokens. These human mesh tokens are then randomly masked, leaving only $M<N$ visible tokens. An embedding layer converts the visible token indices into learned token embeddings, which are subsequently passed to an encoder with multi-head self-attention. The sequence of encoded token embeddings is completed with mask tokens and, during supervised training and inference stages only, it is also concatenated with a sequence of image embeddings (see next paragraph). The complete sequence of tokens is then processed by a decoder with multi-head self-attention, followed by an MLP that outputs a distribution over the $N$ human mesh token indices (in practice, it outputs the logits, which can be normalized using a softmax function). During training and in deterministic inference mode, the $N-M$ human mesh tokens that were originally masked are predicted by taking the argmax, and the Mesh-VQ-VAE decoder is used to reconstruct the canonical mesh $\hat{V}_{c}$. A simple cross-entropy loss is computed from the predicted mesh tokens for both the self-supervised and supervised training stages. In stochastic inference mode, the human mesh tokens are predicted iteratively by sampling the predicted distribution over the mesh token indices and reintroducing a proportion of unmasked tokens at the encoder’s input. Finally, an MLP (not represented in Fig. 2) is trained to predict a global 6D rotation $\hat{R}\in\mathbb{R}^{6}$ and the perspective camera parameters $\hat{\pi}\in\mathbb{R}^{3}$ from image features. These can be used to orient the predicted canonical mesh and to reproject it on the input image.
Image embeddings extractor. An image features extractor computes image features $X\in\mathbb{R}^{WH\times C}$. For fair comparisons with SOTA methods, we rely on an HRNet [79] backbone ($W=H=7$ and $C=720$) pre-trained on a 2D pose estimation task [49]. Additionally, we provide results using a more powerful ViT [17] backbone ($W=12$, $H=16$, $C=1280$). For the multi-output HMR experiment, we adhere to the common practice in the literature and use a ResNet-50 [28] backbone ($W=H=7$, $C=2048$). Prior to being fed to the MEGA decoder, the image features are linearly projected to the Transformer dimension, resulting in $WH$ image embeddings of dimension $D=1024$.
@@ 3.3 Training strategy
Masked generative modeling for HMR. To train MEGA to predict tokenized canonical meshes from images, we extend the pre-training strategy by conditioning the decoder with the image embeddings (see Sec. 3.2 and Fig. 2). The human mesh tokens follow the same masking rate schedule, while the image embeddings remain fully visible. The only supervision for predicting canonical meshes is a cross-entropy loss, as in the pre-training stage. For predicting the rotation and camera parameters, we use the Euclidean norm on the rotation matrix corresponding to the predicted 6D representation and an $L1$ loss on the reprojection of 2D joints extracted from the predicted oriented mesh, using the predicted perspective camera parameters.
3.4 Generation strategy
Deterministic mode. In deterministic inference mode, we predict all tokens in a single forward pass by taking the argmax of the predicted distribution over the human mesh token indices. In this mode, MEGA’s encoder is not used; instead, the decoder is fed with a sequence of $N$ mask tokens, relying entirely on the image representation information. Consequently, the encoder can be discarded, significantly reducing the model size as $B_{e}>B_{d}$. To our knowledge, this is the first work to discard the encoder of an MAE, whereas previous works typically discard the decoder and use the encoder to obtain representations for downstream tasks.
  • The currently visible $n_{t-1}$ tokens are fed to the model.
  • For each of the $P_{t}=N-n_{t-1}$ tokens that are still masked, the model outputs unnormalized probabilities over the indices of the Mesh-VQ-VAE codebook. These probabilities are sampled using the Gumble-max trick [30] to obtain $P_{t}$ candidate tokens to be set visible in the next step. Each candidate token is identified by an index between 1 and $S$ (the codebook size) along with its unnormalized probability.
Given the stochastic nature of this generation mode, we can obtain $Q$ different human mesh predictions for a single image by repeating the above generation process several times. The uncertainty of each vertex can then be computed as its standard deviation over the $Q$ predictions.
4 Experiments
4.1 Experimental setup
Datasets. MEGA is initially pre-trained in a self-supervised manner on a diverse subset of the AMASS dataset [54], focusing on samples with high pose and body shape variety, as detailed in [5]. MEGA is then trained for HMR using a mix of standard image datasets labeled with human meshes, including MSCOCO [49], Human3.6M [31], MPI-INF-3DHP [56] and MPII [1]. We evaluate MEGA and compare it to the SOTA methods on the in-the-wild 3DPW [55] and EMDB [35] datasets. Following recent works [19, 24], we do not finetune MEGA on the 3DPW training set before evaluation. This approach better validates the model’s generalization capacity and allows us to use the same model for all experiments. Unless specified otherwise, other models in the comparison tables use the same training datasets and the same image feature extractor as ours for fair comparisons.
Implementation details. All experiments are conducted using PyTorch [61]. MEGA is pre-trained on human meshes for 500 epochs. For the HMR task, MEGA is first trained on MSCOCO [49] for 100 epochs, followed by training on the mix of datasets described above for 10 epochs. Using 4 NVIDIA A100 GPUs, the entire pre-training and training process takes approximately 3 days.
Metrics. To evaluate HMR methods, we use the widely adopted metrics: per-vertex-error (PVE), mean-per-vertex-error (MPJPE), and the Procrustes-aligned MPJPE (PA-MPJPE). PVE measures the Euclidean distance between the vertices of the predicted mesh and the ground truth. MPJPE assesses the accuracy of body joints extracted from the mesh. PA-MPJPE is similar to MPJPE but includes a Procrustes-alignment, a rigid transformation that minimizes the distance between the predicted and ground truth joints. All three metrics are reported in mm.
Table 1: Evaluation in deterministic mode. We evaluate MEGA on the 3DPW and EMDB datasets and compare it to the SOTA methods using metrics defined in Sec. 4.1 given in mm. $\dagger$ stands for additionally using 2D training data, and $\mp$ for additionally using 2D data and Bedlam [5].
@@ 4.1 Experimental setup
MEGA w/o pre-training (ours) HRNet-w48 84.1 70.5 46.2 113.9 95.9 62.0
MEGA (ours) HRNet-w48 81.6 68.5 44.1 107.9 90.5 58.7

### v3

# MEGA: Masked Generative Autoencoder for Human Mesh Recovery

@@ Abstract
Human Mesh Recovery (HMR) from a single RGB image is a highly ambiguous problem, as an infinite set of 3D interpretations can explain the 2D observation equally well. Nevertheless, most HMR methods overlook this issue and make a single prediction without accounting for this ambiguity. A few approaches generate a distribution of human meshes, enabling the sampling of multiple predictions; however, none of them is competitive with the latest single-output model when making a single prediction. This work proposes a new approach based on masked generative modeling. By tokenizing the human pose and shape, we formulate the HMR task as generating a sequence of discrete tokens conditioned on an input image. We introduce MEGA, a MaskEd Generative Autoencoder trained to recover human meshes from images and partial human mesh token sequences. Given an image, our flexible generation scheme allows us to predict a single human mesh in deterministic mode or to generate multiple human meshes in stochastic mode. Experiments on in-the-wild benchmarks show that MEGA achieves state-of-the-art performance in deterministic and stochastic modes, outperforming single-output and multi-output approaches. Code and trained models will be released upon acceptance.
@@ 1 Introduction
Figure 1: Human mesh recovery from a single image is an ill-posed problem due to depth ambiguity. Probabilistic approaches have aimed to address this by generating multiple predictions, but diversity often sacrifices accuracy. Introducing MEGA, our HMR model based on masked generative modeling achieves state-of-the-art performance on in-the-wild benchmarks in single- and multi-output settings. Given a single image, MEGA can make predictions that all look accurate given the 2D cues but correspond to diverse 3D interpretations.
Perceiving humans from images is a long-standing problem in computer vision, with applications in diverse fields such as sports [23, 83] or e-commerce [55, 93]. Many approaches rely on statistical body models like SMPL [56] for representing humans. Earlier human mesh recovery (HMR) methods recovered the SMPL pose and shape parameters from 2D cues using optimization-based techniques [8, 46]. However, these optimization procedures require good initialization, are time-consuming, and often converge to suboptimal minima. With the advancement of deep learning and the availability of datasets of images with 3D human pose and shape annotations, most approaches have shifted to a regression-based paradigm. Early approaches used neural architectures based on convolutional neural networks (CNNs) and multilayer perceptrons (MLPs) [37, 49]. Recent works have adopted Transformers [80] for extracting image features or making predictions [13, 27]. Despite achieving unprecedented accuracy, state-of-the-art HMR models still have some weaknesses, such as producing unrealistic predictions, especially when dealing with occlusions. To address these issues, recent works have proposed tokenizing the human pose using vector quantized-variational autoencoders (VQ-VAEs) [24, 22]. This approach uses a discrete representation of the human mesh, learned from large-scale motion capture datasets, to confine predictions to the space of anthropomorphic meshes using a dictionary of valid mesh tokens. This tokenization aligns well with Transformer-based architectures, which were initially designed for processing discrete data in natural language processing. Notably, VQ-HPS [24] reframed HMR as a classification task and achieved state-of-the-art results with large and small training datasets, demonstrating the great potential of human mesh tokenized representations in HMR.
While significant progress has been made in HMR, a major issue remains unaddressed in most prior works: a 2D image, especially with occlusions, cannot provide sufficient information to estimate a 3D human mesh with certainty [61, 75] (see  Fig. 1). This limitation causes single-output models to be biased toward most common poses and body shapes [17]. To mitigate this problem, several works have proposed probabilistic approaches that generate multiple predictions from a single image [72, 44]. These approaches have used various families of generative models, ranging from conditional variational autoencoders (CVAEs) [74] to diffusion models [12, 87]. However, this increase in diversity typically comes at the cost of accuracy [73], and none of these multi-output methods are competitive with the latest single-output HMR models when making a single prediction.
@@
In this work, we introduce MEGA, a multi-output HMR approach based on self-supervised learning and masked generative modeling of tokenized human meshes. MEGA relies on the Mesh-VQ-VAE of [24] to encode/decode a 3D human mesh to/from a set of discrete tokens. Our training process unfolds in two steps: (1) Firstly, akin to (vector quantized) masked autoencoders [31, 4, 69, 68], we pre-train MEGA in a self-supervised manner to reconstruct human mesh tokens from partially visible inputs. This leverages an extensive motion capture dataset without needing paired image data, allowing MEGA to learn prior knowledge about 3D humans. (2) Subsequently, for HMR from RGB images, we train MEGA to predict randomly masked human mesh tokens conditioned on image feature embeddings. During inference, we begin with a fully masked sequence of tokens and generate a human mesh conditioned on an input image. We propose two distinct generation modes: (2.a) In deterministic mode, MEGA predicts all tokens in a single forward pass, ensuring speed and accuracy; (2.b) In stochastic mode, the generation process involves iteratively sampling human mesh tokens, enabling MEGA to produce multiple predictions from a single image.
@@
We evaluate MEGA on in-the-wild HMR benchmarks, comparing it to single-output and probabilistic HMR methods. MEGA achieves state-of-the-art (SOTA) performance when predicting a human mesh in a single forward pass in deterministic mode. In stochastic mode, MEGA outperforms both deterministic and probabilistic methods with a single prediction and significantly enhances its performance with an increased number of samples. This mode generates diverse, realistic human meshes, allowing for the proposal of multiple plausible outputs given an image. This allows the user to choose the best-suited solution depending on the use case. For instance, animation designers may want to select the most visually appealing human mesh, while medical practitioners would require high precision and choose the solution that minimizes a reprojection error.
In summary, we make the following key contributions:
@@ 2.1 Human mesh recovery
Single output HMR. Since the release of the HMR [37] model, most approaches for recovering human meshes from images have been regression-based, using neural networks to make predictions directly from the image. These regression-based HMR methods can be categorized into parametric and non-parametric approaches. Parametric methods aim to recover the parameters of the SMPL model [76, 41, 48, 49, 27, 22, 42, 36, 91, 39]. They typically produce realistic predictions; however, some works have argued that the SMPL model parameter space is not the most suitable for predicting human meshes [17, 43, 14], leading to the development of non-parametric approaches. Non-parametric methods predict the coordinates of 3D vertices without relying on a parametric model. While earlier approaches used graph convolutional neural network architectures inspired by the mesh topology [43, 50], recent non-parametric models predominantly employ Transformers [13, 51, 21, 24, 58]. Although non-parametric methods yield accurate results, they can sometimes produce non-anthropomorphic meshes, particularly when training data is scarce [24].
In this work, instead of predicting SMPL model parameters or 3D coordinates, we aim to recover sequences of indices that can be decoded into a human mesh using the Mesh-VQ-VAE from [24]. Thus, the closest methods to ours are TokenHMR [22] and VQ-HPS [24], which rely on a quantized representation of human poses and meshes, respectively. However, the three methods have very distinct ways to address HMR: TokenHMR [22] is a regression method that uses tokens as an intermediate representation; VQ-HPS [24] is a classification method; MEGA is a generation method, which collapses to a classification method in the deterministic mode. Aside from this paradigm difference, MEGA takes partial mesh token sequences as input, while others are fed images. This allows us to first learn an unconditional human mesh generation model only on motion capture data (see Sec. C) before conditioning the generation with images. Ablation studies (see Tab. 1) show that this pre-training without any image conditioning significantly enhances the performance of MEGA.
Multi-output HMR. Estimating a 3D human mesh from a single image is challenging due to the depth ambiguity, especially when the person is partially occluded. Several works have proposed making multiple predictions to account for the ill-posed nature of the problem. Earlier works employed compositional models [35] or mixture density networks [47]. More recent approaches rely on sophisticated probabilistic distributions [72, 71] and generative models, such as CVAE [74], normalizing flows [44, 6, 73], and diffusion models [12, 87]. While these methods can predict diverse plausible solutions, they often face a trade-off between accuracy and diversity [73, 87] and need to make a large number of predictions to be competitive with single-output methods.
@@ 2.2 Self-supervised learning for HMR
While prior works have demonstrated the importance of pre-training for the backbone of HMR models [7, 64], we propose pre-training the generative model on human meshes to leverage extensive motion capture data. Different from prior works in this area [5, 52], our approach based on human mesh tokens does not need any rendering or reprojection, the masking happens in 3D. This pre-training provides us with an unconditional human mesh generative model (see Sec. C), and ablation studies in Tab. 1 show its important contribution to training the HMR model.
@@ 2.3 Masked generative modeling
Masked modeling was introduced in BERT [19] for language modeling and extended to images with the masked autoencoder [31]. This technique trains a model to predict randomly masked tokens in a sequence based on visible tokens. Masked generative modeling builds on this by training a model to generate new samples, starting from a fully masked sequence and iteratively predicting a fixed number of tokens at each step [9, 10, 28].
In this work, we develop the first masked generative modeling approach for HMR. Using a tokenized representation of the human mesh is particularly well-suited for this task, as it allows for straightforward masking and replacement of mesh parts with mask tokens.
@@ 3 MEGA
3.1 Preliminary: Human mesh tokenization
@@
Figure 3: Prediction process iterations. We visualize the predictions for intermediate steps in stochastic mode. All masked tokens are replaced by the first token of the codebook, corresponding to index 0.
@@ 3.2 Model
Overall pipeline. MEGA is a masked generative model based on an encoder-decoder Transformer architecture, illustrated in Fig. 2. During training, the Mesh-VQ-VAE encoder converts a 3D human mesh into a sequence of $N$ tokens. These human mesh tokens are then randomly masked, leaving only $M<N$ visible tokens. An embedding layer converts the visible token indices into learned token embeddings, which are subsequently passed to an encoder with multi-head self-attention. The sequence of encoded token embeddings is completed with mask tokens and, during supervised training and inference stages only, it is also concatenated with a sequence of image embeddings (see next paragraph). The complete sequence of tokens is then processed by a decoder with multi-head self-attention, followed by an MLP that outputs a distribution over the $N$ human mesh token indices (in practice, it outputs the logits, which can be normalized using a softmax function). During training and in deterministic inference mode, the $N-M$ human mesh tokens that were originally masked are predicted by taking the argmax, and the Mesh-VQ-VAE decoder is used to reconstruct the canonical mesh $\hat{V}_{c}$. The cross-entropy loss is computed from the predicted mesh tokens for both the self-supervised and supervised training stages. At test time, the input sequence of mesh tokens is entirely masked (i.e., $M=0$) such that the decoder is provided with $N$ mask tokens along with the conditioning image embeddings. In deterministic inference mode, the $N$ mesh tokens are predicted in a single forward pass. In stochastic inference mode, the human mesh tokens are predicted iteratively by sampling the predicted distribution over the mesh token indices and reintroducing a proportion of unmasked tokens at the encoder’s input. Finally, an MLP (not represented in Fig. 2) is trained to predict a global 6D rotation $\hat{R}\in\mathbb{R}^{6}$ and the perspective camera parameters $\hat{\pi}\in\mathbb{R}^{3}$ from image features. These can be used to orient the predicted canonical mesh and to reproject it on the input image.
Image embeddings extractor. An image features extractor computes image features $X\in\mathbb{R}^{WH\times C}$. For fair comparisons with SOTA methods, we rely on an HRNet [84] backbone ($W=H=7$ and $C=720$) pre-trained on a 2D pose estimation task [53]. Additionally, we provide results using a more powerful ViT [20] backbone ($W=12$, $H=16$, $C=1280$). For the multi-output HMR experiment, we adhere to the common practice in the literature [12, 6, 44] and use a ResNet-50 [30] backbone ($W=H=7$, $C=2048$). Prior to being fed to the MEGA decoder, the image features are linearly projected to the Transformer dimension, resulting in $WH$ image embeddings of dimension $D=1024$.
@@ 3.3 Training strategy
Masked generative modeling for HMR. To train MEGA to predict tokenized canonical meshes from images, we extend the pre-training strategy by conditioning the decoder with the image embeddings (see Sec. 3.2 and Fig. 2). The human mesh tokens follow the same masking rate schedule, while the image embeddings remain fully visible. The only supervision for predicting canonical meshes is a cross-entropy loss, as in the pre-training stage. Compared to prior works in HMR that rely on multiple losses (on 3D keypoints, the 2D reprojection, and the SMPL parameters), this approach is straightforward and does not require hyperparameters tuning. For predicting the rotation and camera parameters, we use the Euclidean distance on the rotation matrix corresponding to the predicted 6D representation and the L1 loss on the reprojection of 2D joints extracted from the predicted oriented mesh, using the predicted perspective camera parameters.
@@
MEGA (ours) HRNet-w48 81.6 68.5 44.1 107.9 90.5 58.7
        linear masking HRNet-w48 86.5 72.6 45.9 118.7 100.1 63.3
        full mask HRNet-w48 81.8 68.5 44.4 110.3 92.7 59.2
        w/o pre-training + full mask HRNet-w48 84.1 70.5 46.2 113.9 95.9 62.0
@@
Table 1: Evaluation in deterministic mode. We evaluate MEGA on the 3DPW and EMDB datasets and compare it to the SOTA methods using metrics defined in Sec. 4.1 given in mm. $\dagger$ stands for additionally using 2D training data, and $\mp$ for additionally using 2D data and BEDLAM [7]. Methods in italic below the row ”MEGA” indicate the results of the ablation study.
3.4 Generation strategy
Deterministic mode. In deterministic inference mode, we predict all tokens in a single forward pass by taking the argmax of the predicted distribution over the human mesh token indices. In this mode, MEGA’s encoder is not used; instead, the decoder is fed with a sequence of $N$ mask tokens, relying entirely on the image representation information. Consequently, the encoder can be discarded, significantly reducing the model size as $B_{e}>B_{d}$. To our knowledge, this is the first work that discards the encoder of an MAE, whereas previous works [31] typically discard the decoder and use the encoder to obtain representations for downstream tasks.
  • The currently visible $n_{t-1}$ tokens are fed to the model.
  • For each of the $P_{t}=N-n_{t-1}$ tokens that are still masked, the model outputs unnormalized probabilities over the indices of the Mesh-VQ-VAE codebook. The corresponding categorical distributions are sampled using the Gumble-max trick [33] to obtain $P_{t}$ candidate tokens to be set visible in the next step. Each candidate token is identified by an index between 1 and $S$ (the codebook size) along with its unnormalized probability.
Given the stochastic nature of this generation mode, we can obtain $Q$ different human mesh predictions for a single image by repeating the above generation process several times. A visualization of the iterative generation process is shown in Fig. 3.
4 Experiments
4.1 Experimental setup
Datasets. MEGA is initially pre-trained in a self-supervised manner on an extensive subset of the AMASS dataset [59], focusing on samples with high pose and body shape variety, as detailed in [7]. MEGA is then trained for HMR using a mix of standard image datasets labeled with pseudo-groundtruth human meshes [49], including MSCOCO [53], Human3.6M [34], MPI-INF-3DHP [60] and MPII [1]. We evaluate MEGA and compare it to the SOTA methods on the in-the-wild 3DPW [81] and EMDB [38] datasets. Following recent works [22, 27, 24], we do not finetune MEGA on the 3DPW training set before evaluation. This approach better validates the model’s generalization capacity and allows us to use the same model for all experiments. Unless specified otherwise, other models in the comparison tables use the same training datasets and the same image feature extractor as ours for fair comparisons.
Implementation details. All experiments are conducted using PyTorch [65]. MEGA is pre-trained on human meshes for 500 epochs. For the HMR task, MEGA is first trained on MSCOCO [53] for 100 epochs, followed by training on the mix of datasets described above for 10 epochs. Using 4 NVIDIA A100 GPUs, the entire pre-training and training process takes about 2.5 days.
Metrics. To evaluate HMR methods, we use the widely adopted metrics: per-vertex-error (PVE), mean-per-joint-error (MPJPE), and the Procrustes-aligned MPJPE (PA-MPJPE). PVE measures the Euclidean distance between the vertices of the predicted mesh and the ground truth. MPJPE assesses the accuracy of body joints extracted from the mesh. PA-MPJPE is similar to MPJPE but includes a Procrustes-alignment, a rigid transformation that minimizes the distance between the predicted and ground truth joints. All three metrics are reported in mm.
3DPW-OCC
Method PVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
ROMP [76] - - 65.9
SPIN [42] 121.6 95.6 60.8
VisDB [89] 110.5 87.3 56.0
3DCrowdNet [15] 103.2 88.6 56.8
SEFD [88] 97.1 83.5 55.0
MEGA (ours) 93.8 79.8 51.5
PARE [41] 107.9 90.5 57.1
ScoreHypo [87] 89.8 73.9 48.7
MEGA (ours) 78.9 66.3 43.7
Table 2: Evaluation on 3DPW-OCC. We evaluate MEGA on an occlusion dataset and compare it to SOTA HMR methods designed to handle occlusions using standard metrics (see Sec. 4.1) in mm. Methods in the top part use a ResNet-50 backbone, while others use HRNet.